OpenStudy (anonymous):

integrate -x(1+2x)^5 dx by parts

hartnn (hartnn):

do we "have to" integrate by parts ? because a simple substitution will make this integral lot easier to solve

OpenStudy (anonymous):

yes the question requires it to be solve by parts.

hartnn (hartnn):

so, do you know integration by parts ? have u tried this ?

OpenStudy (anonymous):

Going with the sub first would probably help make the by-parts-integration less messy.

OpenStudy (anonymous):

i've tried it before but now i forgot from where i should start :( @hartnn

hartnn (hartnn):

do you know the formula for integrating by parts ? (actually there are 2 equivalent formulas, and i wanna know which one do u use)

OpenStudy (anonymous):

integrate udv = uv - integrate vdu

hartnn (hartnn):

so whats your choice of u and dv ?

OpenStudy (anonymous):

idk i tried both i i ciuld find the answers T.T

OpenStudy (anonymous):

*couldnt

OpenStudy (anonymous):

did you get it now ?

OpenStudy (anonymous):

yes thank you so much @sami-21

sam (.sam.):

@sami-21 From \[-\frac{x(1+2x)^6}{12}+\frac{1}{12}\int\limits(1+2x)^6 dx \\ \\ -\frac{x(1+2x)^6}{12}+\frac{1}{12}(\frac{(1+2x)^7}{7(2)})+c \\ \\ -\frac{x(1+2x)^6}{12}+\frac{(1+2x)^7}{168}+c\]

OpenStudy (anonymous):

i missed the 2 ! thnx @.Sam. for clarification.

OpenStudy (anonymous):

yes i got it thank you :)